### v2

# DUDF: Differentiable Unsigned Distance Fields with Hyperbolic Scaling

Abstract
In recent years, there has been a growing interest in training Neural Networks to approximate Unsigned Distance Fields (UDFs) for representing open surfaces in the context of 3D reconstruction. However, UDFs are non-differentiable at the zero level set which leads to significant errors in distances and gradients, generally resulting in fragmented and discontinuous surfaces. In this paper, we propose to learn a hyperbolic scaling of the unsigned distance field, which defines a new Eikonal problem with distinct boundary conditions. This allows our formulation to integrate seamlessly with state-of-the-art continuously differentiable implicit neural representation networks, largely applied in the literature to represent signed distance fields. Our approach not only addresses the challenge of open surface representation but also demonstrates significant improvement in reconstruction quality and training performance. Moreover, the unlocked field’s differentiability allows the accurate computation of essential topological properties such as normal directions and curvatures, pervasive in downstream tasks such as rendering. Through extensive experiments, we validate our approach across various data sets and against competitive baselines. The results demonstrate enhanced accuracy and up to an order of magnitude increase in speed compared to previous methods.
1 Introduction
Surface representation is a fundamental aspect in the field of 3D geometry processing, with explicit methods such as meshes, point clouds, and voxelized representations being traditional choices. Implicit surface representations, on the other hand, have been an integral part of the graphics pipeline for many decades. They encapsulate surfaces as the zero-level set of a function, providing a compact and continuous geometry representation. The novelty in recent years has emerged from parameterizing these implicit functions with Neural Networks (NNs), combining their learning capabilities with the advantages of implicit representations.
Figure 1: Distance field cross-sections comparative. First row displays ground truth unsigned distance fields. Methods based on signed distances, like SIREN [35], mistakenly fill gaps in open surfaces. While methods based on unsigned distances such as CAP-UDF [45] can represent non-watertight surfaces, the learned distance field do not approximate the true function. DUDF’s differentiable formulation successfully overcomes these challenges.
Signed Distance Functions (SDFs) have traditionally been the chosen formulation for implicit surface representation [8, 25] due to their well-defined gradients and the ease they offer for computing constructive solid geometry operations and mesh reconstruction. However, they are inherently limited to closed surfaces (Fig. 1), which poses a significant challenge for representing open surfaces with implicit methods. This limitation arises from the inside/outside sign flip on the zero level set, which is impossible to define for surfaces that do not enclose a volume.
The advent of Unsigned Distance Functions (UDFs) extended representation capabilities to open surfaces. However, this advancement introduced a challenge: the non-differentiable nature of the distance functions at the zero level set. Such non-differentiability leads to inaccuracies in distances and gradients learned by NNs, particularly near the surface where precision is paramount. This paper tackles the challenge of representing open surfaces with neural networks, focusing on a formulation that ensures the surface continuity and smoothness. Our framework (DUDF) combines a hyperbolic scaling of the distance field, together with a new Eikonal problem featuring tailored boundary conditions. This formulation facilitates the training of continuously differentiable implicit neural representation networks, while preserving the essential attributes of the unsigned distance fields. Additionally, the differentiable nature of our representations enables precise calculations of crucial topological characteristics, including curvatures and normals, which are pervasive in downstream tasks such as rendering. In contrast, earlier techniques depended on mesh reconstruction methods like Marching Cubes [19, 14, 46] for rendering purposes.
We evaluate our framework with experiments that rigorously assess performance on a variety of challenging data sets and compare to competitive benchmarks in the field. Our method overcomes several limitations of previous approaches that struggled with the non-differentiability of UDFs. Results demonstrate that our approach not only enhances the reconstruction quality of open surfaces but also accelerates the training process. We believe the dual improvement in precision and efficiency enhance the applicability of neural UDFs representations in geometric processing for real-world scenarios.
2 Related Works
The representation and reconstruction of three-dimensional geometry is a fundamental challenge in the field of computer graphics, vision, and computational geometry. We now briefly discuss the key contributions and methodologies adopted in recent literature to address the main challenges in the field, distinguishing between closed and open surface paradigms.
Closed surfaces. Recent years have seen a surge in the use of Neural Implicit Functions (NIFs) for modeling and reconstructing 3D shapes [39]. These representations are usually implemented with Multi-Layer Perceptrons (MLP) or Convolutional Neural Networks (CNNs), but differ in the learning task [21, 4, 29, 5]. On the one hand, some methods [21, 4, 29, 22, 5] learn an indicator function or binary occupancies which are used to reconstruct the 3D surface. On the other hand, alternative methods focus on estimating the SDF at any point in 3D space [27, 15, 23, 9, 11, 24]. The landscape of NIFs has further expanded with the introduction of novel approaches like implicit moving least-squares surfaces [16], a differentiable Poisson solver [28], a complex Gabor wavelet [32], and a level set alignment loss [20]. Notably, central works in the field [13, 35] approach the problem as finding solutions to the Eikonal equation alongside expressive boundary conditions, building upon the literature on solving PDEs with neural networks [34, 30]. In particular, SIREN [35] achieves remarkable results utilizing periodic activation functions, allowing to successfully control the function’s differential fields. Despite these advancements, a fundamental limitation of current NIF approaches remains in their inability to represent open surfaces, a characteristic often exhibited by real-world objects like scene walls, clothing, or vehicles with inner structures.
Open surfaces. In order to model general non-watertight surfaces, Chibane et al. introduced the idea of learning UDFs, pioneering the handling of open surfaces with neural networks [6]. Following this, several methods have aimed to improve the performance of open-surface representation using neural networks [36, 37, 42, 41, 43, 45]. For instance, GIFS [42] and Neural Vector Fields [41] model the relationships between every two points instead of the relationships between points and surfaces. HSDF separately learns sign and distance fields to handle surfaces with arbitrary topologies[38]. NeuralUDF [18] and NeUDF [17] focus on learning an unsigned distance field by volume rendering for multi-view reconstruction of surfaces with arbitrary topologies. Closer to our work is CAP-UDF [45], that optimizes models on raw point clouds by learning to move 3D point queries until reaching the surface with a field consistency constraint. Despite these developments, since UDFs are not differentiable at the zero level set, the gradient field is ill-defined near the isosurface. This leads to difficulties during training, but specially during surface reconstruction, where gradients are often used to determine surface presence between points [14, 46], compute surface normals, and project points onto the isosurface [6] through gradient descent, among others. Recent work by Zhou et al. aimed to solve this issue by introducing constraints at the zero level set in the form of losses leading to a smoother surface [44]. In contrast, our approach formulates an Eikonal problem which is solved by learning a scaled distance function that remains differentiable close to the surface. We demonstrate that our method not only enhances reconstruction quality, addressing the prevailing challenges associated with the non-differentiability of the unsigned distance, but also improves state-of-the-art in terms of training time performance.
3 Proposed Approach
3.1 Mathematical background
Methods addressing closed surfaces approach the learning of SDFs as finding the solution to a system of Partial Differential Equations (PDEs) governed by the homogeneous Eikonal equation with Dirichlet and Neumann boundary conditions. Formally, given a closed surface $\mathcal{S}$ in $\mathcal{C}\subseteq\mathbb{R}^{3}$ they seek to find a continuous function $f$ that satisfies $\left\|\nabla f\right\|=1$, with boundary conditions $f_{\rvert_{\mathcal{S}}}=0$, $\nabla f_{\rvert_{\mathcal{S}}}=\mathbf{n}_{\mathcal{S}}$. Where $\mathbf{n}_{\mathcal{S}}$ denotes the unitary normal field at the surface. Although SDFs are not differentiable at every point and represent only a weak solution to the Eikonal equation, recent work has demonstrated considerable success in solving this problem with continuously differentiable implicit neural representation networks [35, 13, 26, 7]. These architectures use periodic activations, facilitating smoother optimization processes and improved control over the solution’s gradient field. This is achievable because locations where the signed distance lacks differentiability are distant from the isosurface, hence the approximation tends to be good in a close neighborhood of the zero level set. In particular, these networks enable accurate computation of critical topological properties, including mean and Gaussian curvature [26]. This contrasts with networks featuring piecewise linear activations, such as ReLUs, which have null second-order derivatives [35].
In the context of representing UDFs, these functions cannot be a solution to the Eikonal equation both in the interior and exterior regions of the surface, without losing differentiability at the zero level set. This presents challenges for continuously differentiable implicit neural representation networks in achieving satisfactory outcomes, being that the approximation errors happen at the isosurface where accuracy is paramount. To our knowledge, there has been no successful report of a solution to the Eikonal equation accurately approximating unsigned distance functions.
Figure 2: In (a), view of different distance fields for a 2D slice of a torus. Note the effect of hyperbolic scaling near the isosurface (bottom). In (b), sign and gradient for each distance field $f$. In signed distance, the gradient $\nabla f$ at the isosurface is equal to the surface’s normal field $\mathbf{n}_{\mathcal{S}}$ (top). In unsigned distance, the gradient is undefined at the isosurface due to the change in orientation (middle). Our maximum curvature field vectors $\mathbf{v}_{1}$ align with the surface’s unoriented normals (bottom).
3.2 Problem statement
Our first insight is to redefine the unsigned distance field through the application of a hyperbolic scaling. In light of this understanding, we propose to learn the parameters $\theta$ of a neural network $f_{\theta}$ with periodic activation functions [35] to approximate the function:
$\mathit{t_{\mathcal{S}}}(\mathbf{x})=\mathit{d}_{\mathcal{S}}(\mathbf{x})\tanh(\alpha\mathit{d}_{\mathcal{S}}(\mathbf{x})),$ (1)
where $\mathit{d}_{\mathcal{S}}$ is the unsigned euclidean distance to surface $\mathcal{S}$; and $\alpha$ is a constant value. The function $\mathit{t}_{\mathcal{S}}$ is a differentiable approximation of $\mathit{d}_{\mathcal{S}}$, whose zero level set is surface $\mathcal{S}$. Fig. 2 (a) illustrates hyperbolic scaling’s quadratic smoothing near the isosurface and the linear behavior in the distance. The parameter $\alpha$ controls the distance at which the smoothing occurs.
Our second insight is that the distance scaling enables the application of continuously differentiable implicit neural representation networks to solve an Eikonal equation, while retaining the UDF’s open surface representation capabilities. For this task, we aim to address a heterogeneous Eikonal equation, for which we know $t_{\mathcal{S}}$ (Eq. 1) is a weak solution:
$\left\{\begin{matrix}\left\|\nabla f\right\|=\phi\\ f_{\rvert_{\mathcal{S}}}=0\\ \nabla f_{\rvert_{\mathcal{S}}}=\mathbf{0}\end{matrix}\right.$ (2)
where $\phi$ is defined as the L2 norm of the gradient of Eq. 1, formally:
$\phi(\mathbf{x})=\tanh(\alpha\mathit{d}_{\mathcal{S}}(\mathbf{x}))+\alpha\mathit{d}_{\mathcal{S}}(\mathbf{x})(1-\tanh^{2}(\alpha\mathit{d}_{\mathcal{S}}(\mathbf{x}))).$ (3)
This formulation assumes that the domain boundary is defined by a closed surface $\mathcal{S}$, where Dirichlet and Neumann boundary conditions can be imposed. However, that framework does not directly apply to open surfaces, which do not enclose a well-defined domain and thus cannot support the imposition of boundary constraints. To address this limitation, we reformulate the problem as an initial value problem where the Dirichlet condition on $\mathcal{S}$ prescribes the initial values for $f$, and where the Neumann condition can be omitted given that is already satisfied by the Eikonal equation when $\phi(\mathbf{s})=0$ for $\mathbf{s}\in\mathcal{S}$.
The problem outlined in Eq. 2 does not incorporate the information of the surface normals. This contrasts with prior research focused on closed surfaces, where leveraging such information has led to improved accuracy and enhanced reconstruction quality by further constraining the set of feasible solutions. Our third key insight is that although the gradient of $t_{\mathcal{S}}$ (Eq. 1) becomes null at the isosurface, the direction in which the gradient norm increases most rapidly remains non-null. This direction is indicated by the maximum curvature of $f$, defined as the eigenvector associated with the largest eigenvalue of the Hessian matrix. Note that in this context, the maximum curvature does not refer to the surface’s curvature, but rather to the curvature of the hyperbolic scaled unsigned distance field $f$, where the function’s rate of change varies most rapidly. Formally, $\mathbf{H}_{t_{\mathcal{S}}}$ is the Hessian matrix of $t_{\mathcal{S}}$ which has 3 real eigenvalues $\left|\lambda_{1}\right|\geq\left|\lambda_{2}\right|\geq\left|\lambda_{3}\right|$; let $\mathbf{v_{1}}$ be a unitary eigenvector associated with $\lambda_{1}$ at $\mathbf{s}$, it can be shown that $\mathbf{v_{1}}=\pm\mathbf{n}_{\mathcal{S}}$ for every $\mathbf{s}\in\mathcal{S}$. Therefore, we further condition the solution to Eq. 2 by adding an extra boundary condition enforcing alignment between the unitary directions associated to the maximum curvature field and the unitary normal field of the surface.
3.3 Implicit function learning
Following the aforementioned definitions, our neural networks are trained to minimize the following loss function:
[…]
with $\lambda_{i}$ constant weights controlling the relevance of each term. The term which favors a solution to the Eikonal PDE is defined as:
[…]
The Dirichlet boundary condition loss term controlling function values at the surface is formally defined as:
[…]
Similarly to previous works [26, 7], we extend this conditioning to points far from the surface ${\mathcal{S}}$. This is achieved by computing an approximation of the function $t_{\mathcal{S}}$ (Eq. 1) based on nearest neighbors. Neumann’s boundary condition is expressed in the loss term:
[…]
Finally, we ensure the second-order boundary condition by aligning the directions of maximum curvature with the surface normal through the computation of the following integral:
[…]
where $\mathbf{v_{1}}(\cdot)$ is the unitary eigenvector associated with the largest eigenvalue of $\mathbf{H}_{f_{\theta}}(\mathbf{x})$, the Hessian matrix of $f_{\theta}$ at $\mathbf{x}$. In practice, we approximate all integrals discretely using a data set comprised of tuples $\{(\mathbf{x}_{i},\mathit{d}_{\mathcal{S}}(\mathbf{x}_{i}),\mathbf{n}_{\mathcal{S}}(\mathbf{x}_{i}))\}_{i}$. At every training iteration, the gradient and Hessian of the neural network are computed through automatic differentiation.
3.4 Isosurface refinement
In pursuit of solutions that exhibit a consistent, near-zero value at the isosurface —essential for smooth reconstructions— we further fine-tune our networks in a second optimization step. Formally, we minimize the following loss function:
$\mathcal{L}_{Refinement}=\lambda_{\mu}\rvert\,\mu({f_{\theta}}_{\rvert_{\mathcal{S}}})\,\rvert+\lambda_{\sigma}\,\sigma({f_{\theta}}_{\rvert_{\mathcal{S}}}),$ (9)
where $\mu(\cdot)$ is the mean value and $\sigma(\cdot)$ the standard deviation. By minimizing the mean and variance of the learned function values at the isosurface, we ensure they are as close to zero as possible. This refinement process effectively reduces the oscillations and deviations that can occur at the critical boundary, thereby enhancing the overall quality of the reconstructed surfaces.
3.5 Normals and curvature computation
A significant advantage of our formulation over prior efforts [6, 46] is the fact that gradient fields and higher-order derivatives are well-defined in the vicinity of the isosurface, allowing the direct computation of topological properties during inference and reconstruction. This means that we are able to render unsigned distance fields using standard algorithms such as sphere tracing. However, directly using the gradients of $f_{\theta}$ at the isosurface as normals for rendering purposes can be unreliable since their norm is close to 0. Hence, for a given surface value threshold $\epsilon$, when a point $\mathbf{s}$ is identified such that $f_{\theta}(\mathbf{s})<\epsilon$, we determine the normal direction to the surface by calculating the unit eigenvectors associated with the maximum eigenvalue of $\mathbf{H}_{f_{\theta}}(\mathbf{s})$. Note that there are two unitary eigenvectors per surface point, $\mathbf{v_{1}}$ and $-\mathbf{v_{1}}$; however this is usually the desired behavior since open surfaces may not be orientable, such as the möbius strip. To address this ambiguity in the sign we consider the position of the camera when rendering. Additionally, other relevant topological properties such as the mean and Gaussian curvature are also available by means of the maximum unit eigenvector field. See Sec. 4.6 for more details.
4 Results and evaluation
4.1 Experimental setup
We conducted a series of experiments to assess the performance of our method. To target a broad amount of surfaces, we experimented on three well-known data sets: ShapeNet cars [3], Multi-Garment [2], and DeepFashion [48]. For each data set we trained individual networks on 30 randomly selected oriented point clouds. For the Deep Fashion and Multi-Garment dataset, we sampled 100,000 surface points, while for ShapeNet cars we doubled the sample size to account for the higher degree of complexity. On each training iteration, the loss was computed on a subset of 30,000 points, equally distributed into three distinct groups: surface points (1), far domain points (2), and near domain points (3). Surface points (1) were randomly selected from the ground truth point cloud. Far points (2) were uniformly generated within the function’s domain: a cube with side length 2 and centered at the origin. All shapes were uniformly scaled to fit within this volume before sampling. The distance to the point clouds was approximated using a tree-based search algorithm [47]. Finally, since we found that a biased near-the-surface sampling improved accuracy, we used near domain points (3) during training. We constructed this set by randomly displacing the sampled surface points in their normal direction. Displacements were sampled from a normal distribution centered at zero with a standard deviation of $0.01$. The corresponding unsigned distance was approximated with the distance to the undisplaced surface point. This was experimentally found to be sufficiently accurate and fast.
4.2 Network architecture
Our network takes spatial coordinates $x,y,z$ as input, and outputs the predicted $t_{\mathcal{S}}$ value from Eq. 1. The architecture consists of an 8-layer 256-units MLP with sine activations. The same number of layers and parameter count were used for baseline methods. Loss weights were experimentally set to $\lambda_{e},\lambda_{d},\lambda_{n}=1e^{4}$, $\lambda_{g}=1e^{3}$ and $\lambda_{\mu},\lambda_{\sigma}=1e^{5}$. Models were implemented in Python using Pytorch library, and distance metrics were computed using Pytorch3D [31]. We trained each network for 3,000 iterations, the first third of them utilized a learning rate of $1\mathrm{e}^{-4}$, in the following third we lowered it to $1\mathrm{e}^{-5}$, and in the final iterations (only the refinement loss active) we set the learning rate to $1\mathrm{e}^{-7}$ with a cosine decay. These final small learning rates allowed the refinement of the isosurface without losing the gained accuracy in the rest of the field. Regarding $\alpha$ (Eq. 1), we experimentally found that choosing a value of 100 kept a good balance between training accuracy and reconstruction quality. Experiments were run on Ubuntu, an Intel(R) Xeon(R) Silver 4310 CPU, 256Gb of RAM, and 2 Nvidia A100 graphics cards of 80Gb VRAM.
4.3 Mesh reconstruction
For a thorough comparison of our method against existing techniques, we undertake mesh reconstruction using two gradient-based Marching Cubes algorithms, referred here as MC1 [46] and MC2 [14]. These methods perform linear interpolation between grid values for accurate placement of triangle vertices. Given this requirement and considering that the function $t_{\mathcal{S}}$, as described in Equation 1, exhibits quadratic behavior near the isosurface, we retrieve the true unsigned distance by first dividing by $\alpha$ and then applying the square root. We present results using the Chamfer distance (L1,L2), pervasive in related literature. Additionally, we compute the Normal Consistency metric (NC) as the mean value of the absolute cosine similarities between each point to its closest neighbor in the other point cloud.
4.4 Surface reconstruction
4.4.1 Closed shapes
Method time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$
DeepSDF [27] $\mathbf{94}$ $19.40$ $0.206$
SIREN [35] $379$ $15.40$ $0.171$
CAP-UDF [46] $1080$ $\mathbf{9.48}$ $0.030$
Ours (MC1 [46]) $319$ $\mathbf{9.48}$ $\mathbf{0.028}$
Ours (MC2 [14]) $9.49$ $\mathbf{0.028}$
Table 1: Training time, L1, and L2 mean Chamfer distances (×$\times 10^{3}$) for the closed ShapeNet cars data set. DeepSDF and SIREN were trained on signed distances, while CAP-UDF and our method were trained on unsigned distances.
We benchmarked DUDF against state-of-the-art methods for representing closed surfaces, where signed distances are well-defined. Given that our primary focus is not on closed surface representation, we limited this comparison to the ShapeNet car dataset [3], modifying the meshes for closure and omitting internal structures [40]. The comparative results are detailed in Table 1.
While DeepSDF [27] excels in speed, it falls short in reconstruction quality. Being primarily designed for shape generation, DeepSDF is outperformed by methods tailored for accurate single shape representation.
Figure 3: Comparisons on DeepFashion [48] (top row), Multi-Garment [2] (middle row), and ShapeNet cars [3] (bottom row) data sets. DUDF preserves fine details and accurately represents complex geometries without closing holes, outperforming SIREN (SDF), which tends to smooth and round models, and CAP-UDF, which captures sharp features but often closes open surfaces. Reconstructions for CAP-UDF and DUDF performed with MC1 [46].
Despite SIREN [35] not reporting performance on learning distance fields, we included it in our comparison for the sake of fairness and comprehensiveness, given that our method’s backbone is built upon its architecture. When tested on close surfaces, SIREN shows improved shape representation with enough detail and smoothness. However, since its zero level set satisfies the implicit function theorem, it cannot properly represent shapes with right angles and thin structures. In contrast, CAP-UDF [46] can better approximate these sharp features because it does not rely on continuously differentiable NNs, thereby achieving higher accuracy. However, the training time is high and the reliance on gradient based Marching Cubes methods for surface reconstruction tends to create surfaces with small holes and discontinuities which hinders the visual quality in some cases. Our approach captures a differentiable function’s zero level set like SIREN, but uniquely bypasses the implicit function theorem by having null gradients at the isosurface. This allows our method to effectively represent sharp features and complex geometries, leading to improved precision in surface reconstruction, particularly in scenarios where surfaces are not inherently smooth.
4.4.2 Open surfaces
When evaluating open-surface representation, we primarily compare our method against CAP-UDF [46]. This approach was shown to be more effective than earlier techniques (i.e. NDF [6]), that required a complex process for creating a dense point cloud and reconstructing a triangulated surface using the Ball Pivoting algorithm [1]. This often leads to poor quality meshes with holes and irregularities. Additionally, a recent method by Zhou et al. [44] has introduced constraints on the zero level set, which the authors claim it leads to better reconstructions. However, the absence of available implementation code in their official repository precluded a fair comparison with our method. Therefore, we concentrate our analysis on the most recent and available methodologies.
Method
SIREN [35]
CAP-UDF [46]
Ours (MC1 [46])
Ours (MC2 [14])
DeepFashion [48]
time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$ NC $\downarrow$
$376$ $27.3$ $1.980$ $0.107$
$1390$ $18.1$ $1.110$ $0.080$
$\mathbf{326}$ $\mathbf{9.01}$ $\mathbf{0.025}$ $0.024$
$9.14$ $0.027$ $\mathbf{0.020}$
Multi-Garment [2]
time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$ NC $\downarrow$
$374$ $40.5$ $8.810$ $0.094$
$1440$ $18.5$ $1.190$ $0.083$
$\mathbf{318}$ $\mathbf{8.70}$ $\mathbf{0.024}$ $0.026$
$8.82$ $0.026$ $\mathbf{0.021}$
ShapeNet cars [3]
time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$ NC $\downarrow$
$751$ $16.9$ $0.170$ $\mathbf{0.240}$
$1040$ $\mathbf{10.9}$ $0.071$ $0.288$
$\mathbf{317}$ $12.3$ $\mathbf{0.057}$ $0.387$
$13.7$ $0.081$ $0.304$
Table 2: Training time, L1 and L2 mean Chamfer distances (×$\times 10^{3}$), and Normal Consistency (NC) for the evaluated open surface data sets. SIREN was trained as described in the original paper, while CAP-UDF and our method were trained on unsigned distances.
Qualitative comparisons can be observed in Fig. 3 and a full quantitative analysis is presented in Table 2. On the one hand, our method demonstrates a significant improvement in efficiency, consuming up to an order of magnitude less computational time than CAP-UDF, while also showing enhanced performance across all three data sets. This advantage is particularly evident in the DeepFashion and Multi-Garment data sets, where CAP-UDF tends to inaccurately close openings (like those at the ends of sleeves in clothing). On the other hand, SIREN creates a negative distance valued shell around the isosurface. This causes to either close every hole or a thickened surface enclosing the ground-truth level set.
In addition to these experiments, we compare our method with CAP-UDF in the context of sphere tracing rendering. This comparison is presented in Fig. 4, where we showcase two rendering examples. Since learned unsigned distance fields in CAP-UDF do not grow linearly away from the surface, the marching steps in sphere tracing often fail to accurately intersect the surface, leading to undesirable visual artifacts. Even when hitting the surface, gradients might be undefined, leading to noisy images. In contrast, our proposed distance function does not exhibit these problems, offering a more robust framework for direct rendering. Furthermore, our ability to compute normals using the maximum curvature field is useful for shading purposes, enhancing the rendering quality without undergoing an intermediate 3D mesh reconstruction step.
4.5 Ablation study
Figure 4: Rendering examples using sphere tracing. CAP-UDF struggles with non-linear growth of their unsigned distance fields, causing the sphere tracing marching step to miss the surface. Conversely, our method demonstrates precision in direct rendering scenarios.
We now dissect the impact of individual loss components on the network’s accuracy through an ablation study. The methodology involves selectively deactivating terms in our loss function (Eq. 4) to isolate their contributions. We adopted the experimental framework from the prior section, training on 30 meshes from the DeepFashion dataset [48]. We report metrics for reconstructions performed using MC2 [14], yielding results analogous to those obtained with MC1 [46].
The findings presented in Table 3 underscore the significance of the Eikonal equation (Eq. 5) in achieving superior normal consistency ($\lambda_{g}$). Moreover, the refinement step during training ($\lambda_{\mu},\lambda_{\sigma}$) is an important element for normal consistency improvement, offering substantial increased performance for a small additional computational cost. Although $\mathcal{L}_{MCurv}$ (Eq. 8) does not offer substantial quantitative benefits ($\lambda_{g}$), we found that this loss term becomes very relevant qualitatively when rendering through direct methods (see Supp. Material for more details). Additionally, we contrasted the effects of approximating the ground truth distance function $d_{\mathcal{S}}$ (with its distinct Eikonal problem), using the same network architecture and sampling scheme, with sine and ReLU activations. The findings from this study show that such an approach leads to suboptimal outcomes, thereby underscoring the effectiveness of $t_{\mathcal{S}}$ and the derived Eikonal problem proposed.
Finally, we ablated parameter $\alpha$ (see Supp. Material). As $\alpha$ gets larger, function $t_{\mathcal{S}}$ closely approximates $d_{\mathcal{S}}$, increasing the reconstruction error (probably due to the non-differentiability at the isosurface). Smaller $\alpha$ values enlarge the quadratic strip near the isosurface, which is harder to supervise effectively and hinders the performance of MC.
Method time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$ NC $\downarrow$
Baseline $326$ $\mathbf{9.14}$ $\mathbf{0.027}$ $\mathbf{0.020}$
$\lambda_{e}=0$ $312$ $9.43$ $0.028$ $0.033$
$\lambda_{g}=0$ $150$ $9.16$ $\mathbf{0.027}$ $0.021$
$\lambda_{e},\lambda_{g}=0$ $\mathbf{109}$ $9.44$ $\mathbf{0.027}$ $0.035$
$\lambda_{\mu},\lambda_{\sigma}=0$ $302$ $9.24$ $\mathbf{0.027}$ $0.031$
$d_{\mathcal{S}}$ (sine) $150$ $31.2$ $0.830$ $0.057$
$d_{\mathcal{S}}$ ($ReLU$) $145$ $47.5$ $2.500$ $0.149$
Table 3: Quantitative impact of each loss component on the reconstruction accuracy of our network. We report training time, L1, and L2 mean Chamfer distances ×$\times 10^{3}$, and Normal Consistency (NC). Last two rows correspond to approximating the ground truth distance function $d_{\mathcal{S}}$ instead of $t_{\mathcal{S}}$.
4.6 Computing mean and Gaussian curvature
[…]
$H(\mathbf{s})=\frac{1}{2}\nabla\cdot\mathbf{n}_{\mathcal{S}}(\mathbf{s}).$ (10)
Additionally, the Gaussian curvature $K$ can be computed using the determinant of a matrix composed of the Jacobian matrix $J_{\mathbf{n}_{\mathcal{S}}}(\mathbf{s})$ of the unit normal field ($\mathbf{n}_{\mathcal{S}}$) along with the normal field itself. Formally, we compute:
$K(\mathbf{s})=-\det\left[\begin{array}[]{c|c}J_{\mathbf{n}_{\mathcal{S}}}(\mathbf{s})&\mathbf{n}_{\mathcal{S}}(\mathbf{s})\\ \hline\cr\mathbf{n}_{\mathcal{S}}(\mathbf{s})^{t}&0\end{array}\right].$ (11)
In Fig. 5 we show mean and Gaussian curvatures for open and closed surfaces computed with our method.
Figure 5: Mean and Gaussian curvatures computed with our method for closed surfaces (left) and open surfaces (right).
5 Limitations and future work
Our method encounters certain limitations. Primarily, by approximating the function $t_{\mathcal{S}}$ (Eq. 1), the Eikonal equation requires supervision with ground truth unsigned distances, in contrast to signed distance approaches where it is constant everywhere. In addition, the quadratic behavior of the function near the isosurface produces a wider near-zero strip which can lead to slightly inflated representations. This effect can be lessened by computing square roots, at the expense of an increased computational cost. Besides hyperbolic scaling, we experimented with the square distance and SmoothL1. The former requires direct rendering to compute square roots at every marching step and SmoothL1 does not yield the true distance for points far from the isosurface when $\delta$ is nonzero. This is not the case for our hyperbolic function, that satisfies $t_{\mathcal{S}}=d_{\mathcal{S}}$ for distant domain points. Future research can find other suited functions for distance field learning. Secondly, we observed that surfaces extracted using gradient-based Marching Cubes methods can sometimes be less smooth than directly rendering the function using sphere tracing. This is likely due to the current reconstruction algorithm’s sensitivity to noise, an issue that might be addressed by employing a tangency-aware surface reconstruction strategy [33]. Additionally, computing the maximum curvature direction field is computationally expensive, requiring three network passes and the diagonalization of the Hessian matrix. Future network’s designs could reduce the number of parameters, thus lessening the computational burden associated with computing topological features.
6 Conclusion
In this work we introduced Differentiable Unsigned Distance Fields (DUDF) with Hyperbolic Scaling, a novel approach that addresses inherent limitations of traditional UDFs in representing open surfaces. By applying a hyperbolic transformation to the distance field, we define a new variant of the Eikonal problem, tailored with unique boundary conditions. This adaptation enables our model to capitalize on the robust framework of continuously differentiable neural networks, thus enhancing reconstruction quality and training efficiency. The conducted experiments provide evidence that our approach can lead to improvements in reconstruction quality and computational performance when compared to several state-of-the-art methods. Moreover, the ability to accurately calculate topological features such as normals and curvatures is a notable benefit of our model, offering additional utility in geometric processing tasks and rendering. While our results are promising, showing better accuracy and efficiency in most cases, we acknowledge the need for further research and development. The potential applications of DUDF in 3D geometry processing are broad, and we are optimistic about its utility in practical scenarios.
7 Aknowledgments
This project was supported by grants from Salesforce, USA (EinsteinAI 2020), National Scientific and Technical Research Council (CONICET), (PIP 2021-2023 GI -11220200102981CO), and Universidad Torcuato Di Tella, Argentina. We also thank Facundo Fainstein and Tomas Delgado for valuable discussions on the Eikonal problem.